OpenStudy (anonymous):

If f(x) = x minus 3, all over 3, then what is f-1(x) ?

OpenStudy (p0sitr0n):

f(x) = (x-3) /3 you swithc y and x, so x = (y-3)/3 isolate y 3x+3 = y